To use the simplex method to solve the following LLP :

Maximise z = 4x₁ + 3x₂

Subject to :

2x₁ + x₂ <= 1000

x₁ + x₂ <= 800

x₁ <= 400

x₂ <= 700

x₁ , x₂ >= 0


1
Expert's answer
2021-05-10T18:21:29-0400

Max "Z=0.4x_1+0.3x_2"


subject to-

"2x_1+x_2\le 1000\\

x_1+x_2\le 800\\

x_1\le400\\

x_2\le 700\\

x_1,x_2\ge 0"

After introducing slack variables -


"Max Z=0.4x_1+0.3x_2+0s_1+0s_2+0s_3+0s_4"


subject to-


"2x_1+x_2+s_1=1000\\

x_1+x_2+s_2=800\\

x_1+s_3=400\\

x_2+s_4=700\\"

and "s_2,s_2,s_3,s_4,x_1,x_2\ge 0"




Negative minimum "Z_j-C_j" is -0.4 and its column index is 1. So the next value is x_1


Minimum ration is 400 and its row index is 3. so the leaving basis variable is "s_3."

The pivot element is 1. So entering "x_1" , departing s_3 and key element =1



Since all "Z_j-C_j\ge 0"

Hence ,Optimal solution is arrived with value of "x_1=200,x_2=600"


Max "Z=260"